OpenStudy (anonymous):

Which inequality will have a solid boundary line and a shaded region above its graph? x − y ≥ 3 2x − 3y ≤ 3 3y − x < 2 2x + y < 7

pooja195 (pooja195):

\[\huge~\rm~< = dotted ~~~line\] \[\huge~\rm~\le=Solid~~line\]

pooja195 (pooja195):

Which options can we elminate ?

OpenStudy (anonymous):

the bottom two

OpenStudy (mathstudent55):

\(y \lt mx + b\) means shade below dashed line \(y \gt mx + b\) means shade above dashed line \(y \le mx + b\) means shade below solid line \(y \ge mx + b\) means shade above solid line Solve each inequality for y, and compare with 4 statements above. Remember that in solving an inequality, if you multiply or divide both sides by a negative number, the inequality sign changes direction.